Simplify (w^5y^4)^3
Found 2 solutions by mukhopadhyay, Prithwis:
Answer by mukhopadhyay(490)

(w^5y^4)^3
=(w^5)^3(y^4)^3
=w^15y^12

Answer by Prithwis(166)

(w^5y^4)^3
=(w^5)^3(y^4)^3
=w^15y^12 And that is your answer